40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Aug 5 13:28:43 2018, gautam, Update, CDS, c1lsc flaky 
    Reply  Mon Aug 6 00:26:21 2018, gautam, Update, CDS, More CDS woes CDSoverview.png
       Reply  Mon Aug 6 14:38:38 2018, gautam, Update, CDS, More CDS woes CDScrash.pngMC1failures.png
          Reply  Mon Aug 6 19:49:09 2018, gautam, Update, CDS, More CDS woes IMG_7106.JPG
             Reply  Tue Aug 7 11:30:46 2018, gautam, Update, CDS, More CDS woes 
                Reply  Tue Aug 7 22:28:23 2018, gautam, Update, CDS, More CDS woes 
                   Reply  Wed Aug 8 23:03:42 2018, gautam, Update, CDS, c1lsc model started 
                      Reply  Thu Aug 9 12:31:13 2018, gautam, Update, CDS, CDS status update 
                         Reply  Wed Aug 15 21:27:47 2018, gautam, Update, CDS, CDS status update 
                            Reply  Tue Sep 4 10:14:11 2018, gautam, Update, CDS, CDS status update 
                               Reply  Wed Sep 5 10:59:23 2018, wgautam, Update, CDS, CDS status update 
                                  Reply  Thu Sep 6 14:21:26 2018, gautam, Update, CDS, ADC replacement in c1lsc expansion chassis CDSoverview.png
                                     Reply  Fri Sep 7 12:35:14 2018, gautam, Update, CDS, ADC replacement in c1lsc expansion chassis Screenshot_from_2018-09-07_12-34-52.png
                                        Reply  Mon Sep 10 12:44:48 2018, Jon, Update, CDS, ADC replacement in c1lsc expansion chassis 
                                     Reply  Thu Sep 20 11:29:04 2018, gautam, Update, CDS, New PCIe fiber housed PCIeFiberSwap.pngPCIeFiberSwap_FBrebooted.png
                                        Reply  Thu Sep 20 16:19:04 2018, gautam, Update, CDS, New PCIe fiber install postponed to tomorrow 
                                           Reply  Fri Sep 21 16:46:38 2018, gautam, Update, CDS, New PCIe fiber installed and routed PCIeFiber.pngIMG_5878.JPG
Message ID: 14203     Entry time: Thu Sep 20 16:19:04 2018     In reply to: 14202     Reply to this: 14206
Author: gautam 
Type: Update 
Category: CDS 
Subject: New PCIe fiber install postponed to tomorrow 

[steve, gautam]

This didn't go as smoothly as planned. While there were no issues with the new fiber over the ~3 hours that I left it plugged in, I didn't realize the fiber has distinct ends for the "HOST" and "TARGET" (-5 points to me I guess). So while we had plugged in the ends correctly (by accident) for the pre-lunch test, while routing the fiber on the overhead cable tray, we switched the ends (because the "HOST" end of the cable is close to the reel and we felt it would be easier to do the routing the other way. 

Anyway, we will fix this tomorrow. For now, the old fiber was re-connected, and the models are running. IMC is locked.

Quote:

Pictures + more procedural details + proper routing of the protected fiber along cable trays after lunch. If this doesn't help the stability problem, we are out of ideas again, so fingers crossed...

ELOG V3.1.3-